FOURIER TRANSFORM MICROWAVE SPECTROSCOPY OF THE CONFORMERS OF PERFLUOROBU-TANE

\underline{MICHAELEEN R. MUNROW}, S. RANGANATHAN, STEWART E. NOVICK, Department of Chemistry, Wesleyan University, Middletown, CT 06459; DEAN ANTIC, JOSEF MICHL, Department of Chemistry and Biochemistry, University of Colorado, Boulder, CO 80309.

Nitrogen matrix IR spectroscopy and ab initio calculations of three distinct conformations of perfluoro-n-butane, n-C_{4}F_{10}, have recently been reported^{a}. We have measured the Fourier Transform microwave spectrum of the gauche conformer of perfluoro-n-butane and determined the C-C-C-C dihedral angle to be ∼ 51^{∘}. This is in good agreement with the ab initio calculations of Albinsson and Michl that predicted dihedral angles of 54^{∘}, 95^{∘}, and 165^{∘}, for the gauche, ortho, and anti conformations, respectively. The search for the other conformations continues. In addition, the spectrum of the symmetric top isomer, perfluoroisobutane, has been observed.

aB. Albinsson and J. Michl, J. Phys. Chem. 100, 3418 (1996)